lee melts, determine the final temperature; if it does not, determine how much ice melts. P6.8 A single pane glass window of 5 mm thickness and dimensions of 3ft x 2ft is positioned vertically. If the inside temperature is 74 ℉ and the outside temperature is 88 °F, determine how much heat enters the house through the window in 1 hour P6.9 A 600g bottle of water at 297 K is placed in a refrigerator which is able to extract 30 Infheatp6.8
Answer #1

p6.8

Area=0.9144*0.6096

Power=kAT/l

=0.9*0.9144*0.6096*7.77/(5*10^-3)

Power=779.605 W

Energy=3600*779.605

Energy=2.8*10^6 J